# AI Fact-Checking in the Wild: A Field Evaluation of LLM-Written Community Notes on X

### 03 Apr 2026 v1

@@ Introduction
Results
Platform Dynamics Complicate Note-Level Comparisons
@@ Results
LLM Notes Receive More Positive Ratings Across the Ideological Spectrum
@@
To examine whether rating patterns varied across the political spectrum, we stratified raters into three ideology groups based on their rater factor: left (factor <0.15<-0.15), neutral ($0.15\leq$ factor $\leq 0.15$), and right (factor >0.15>0.15). For each note and ideology group, we computed the percentage of ratings marked helpful and not helpful, then aggregated across notes to obtain mean percentages with 95% confidence intervals. Results are presented in Figure 1. LLM notes have higher avg. % helpful ratings and lower avg. % unhelpful ratings than human notes across all rater ideology groups. The largest difference appears among neutral raters, followed by left-leaning raters, and then right-leaning raters.
To quantify differences in individual ratings between LLM and human notes, we estimated a linear mixed effects model predicting individual rating scores from LLM writer authorship, rater political leaning, and their interactions. This analysis mirrors the “bridging” principle of Community Notes algorithm by modeling how the LLM advantage varies as a function of rater ideology, and the extremity of their ideology. Community Notes allows users to rate each note as Helpful, Somewhat Helpful, or Not Helpful; following the platform convention, we coded these responses numerically as 1.0, 0.5, and 0.0, respectively.
Specifically, we use the following primary specification:
$\text{rating\_score}\sim\text{AI}\times\text{coreRaterFactor1}+\text{AI}\times(\text{coreRaterFactor1})^{2}+(1\mid\text{noteId})+(1\mid\text{raterId})$ (1)
@@
Note + Rater RE Tweet + Rater RE OLS Clustered Note + Rater RE (Group) (Intercept) 0.7850.785*** 0.842\$0.842\$*** 0.8260.826*** 0.7850.785*** ($0.006$) (0.005)(\$0.005\$) ($0.003$) ($0.007$) AI 0.1040.104*** 0.066\$0.066\$*** 0.0820.082*** 0.0980.098*** ($0.009$) (0.004)(\$0.004\$) ($0.004$) ($0.010$) coreRaterFactor1 0.015-0.015*** 0.007-0.007 0.005-0.005 ($0.004$) (0.004)(\$0.004\$) ($0.003$) coreRaterFactor12 0.175-0.175*** 0.195\$-0.195\$*** 0.291-0.291*** ($0.010$) (0.010)(\$0.010\$) ($0.009$) AI:coreRaterFactor1 0.087-0.087*** 0.112\$-0.112\$*** 0.120-0.120*** ($0.005$) (0.005)(\$0.005\$) ($0.005$) AI:coreRaterFactor12 0.190-0.190*** 0.169\$-0.169\$*** 0.181-0.181*** ($0.013$) (0.013)(\$0.013\$) ($0.014$) left-leaning rater 0.031-0.031*** ($0.005$) right-leaning rater 0.059-0.059*** ($0.005$) AI:left-leaning rater 0.003-0.003 ($0.007$) AI:right-leaning rater 0.086-0.086*** ($0.007$) SD (Intercept raterParticipantId) $0.160$ $0.160$ $0.164$ SD (Observations) $0.304$ $0.335$ $0.304$ SD (Intercept noteId) $0.196$ $0.197$ SD (Intercept tweetId) $0.142$ Num.Obs. 108 169108\,169 108 169108\,169 108 169108\,169 108 169108\,169
@@
Table 1: Rating-Level linear mixed effects model results. To assess robustness, we estimate three additional specifications. Model 2 replaces note random intercepts with tweet random intercepts — a coarser grouping that absorbs tweet-level confounders while still identifying the AI advantage partly from within-tweet variation. Model 3 is an OLS regression with standard errors clustered by note. Model 4 replaces the continuous coreRaterFactor1 with categorical rater group (left-leaning, neutral, right-leaning; reference: neutral). The estimates of AI advantage stay consistent across all specifications.
@@
Equal-Exposure Note-Level Analysis
@@
LLM and Human Notes Differ in Writing and Sources
LLM and human notes differed in both writing and source citation behaviors. LLM notes were much longer than human notes (mean 35.8 words vs. 26.9 words, $t=22.190$, $p<0.001$). The LLM writer also cited more URLs on average (mean: 1.51 vs. 1.23 URLs; $t=8.67$, $p<0.001$). Analysis of cited domains reveals distinct sourcing patterns (Tables 2 and 3). LLM notes most frequently referenced mainstream news outlets and social media platforms, including reuters.com (7.7% of LLM notes), en.wikipedia.org (7.2%), instagram.com (5.8%), youtube.com (5.3%), x.com (4.2%), bbc.com (4.2%), snopes.com (3.9%), cnn.com (2.4%), facebook.com (2.0%), and yahoo.com (2.0%). In contrast, human notes most commonly cited x.com (18.7%), nearly four times the rate observed in LLM notes (4.2%). Wikipedia appeared at similar rates in both groups (7.6% human vs. 7.2% LLM), but other mainstream news sources appeared less frequently in human notes: reuters.com (1.6% vs. 7.7%), bbc.com (1.1% vs. 4.2%), and snopes.com (0.7% vs. 3.9%). This suggests that the LLM writer relies more heavily on traditional authoritative sources, whereas human writers more frequently reference platform-native content and social media posts. We note that this pattern characterizes our specific implementation.
@@
Table 2: Top 10 domains in LLM notes
@@
Table 3: Top 10 domains in human notes
Discussion
@@
Data and Methods
Dataset
@@
Analysis and multiple testing strategy
@@
Note Writing Pipeline
@@
AI Usage Statement
@@ Appendix B Full-Sample Note-Level Outcomes
% Helpful, mean [95% CI] % Unhelpful, mean [95% CI]
Rater Ideology LLM Human LLM Human
Left 79.55 [77.83, 81.27] 68.10 [65.95, 70.24] 17.98 [16.33, 19.63] 28.72 [26.61, 30.83]
     ($n$: 1,300 / 1,183)
Neutral 89.58 [88.32, 90.84] 76.83 [74.86, 78.79] 07.60 [6.52, 8.69] 18.40 [16.56, 20.25]
     ($n$: 1,174 / 1,049)
Right 73.13 [71.32, 74.93] 67.46 [65.52, 69.41] 23.65 [21.92, 25.38] 28.58 [26.68, 30.48]
     ($n$: 1,300 / 1,196)
@@
Metric LLM Human Test AI coef. Statistic $p$
NN (total notes) 1,614 1,332
CRH rate 13.07% 18.02% LMM 0.058-0.058 z=4.56z=-4.56 <0.001<0.001
CRNH rate 1.12% 4.13% LMM 0.027-0.027 z=4.71z=-4.71 <0.001<0.001
Helpfulness score 0.25 0.24 LMM $0.007$ $z=1.14$ 0.253
     (NN with scores: 1,243 / 1,130)
Num. ratings, median 22 51 Mann–Whitney $U$ U=740,798U=740{,}798 <0.001<0.001
Time to CRH, hrs, median 7.38 5.90 Mann–Whitney $U$ U=27,979U=27{,}979 0.054
     (NN CRH notes: 211 / 240)
@@ Appendix C Robustness Checks for Note-Level Outcomes
Metric LLM Human Test AI coef. Statistic $p$
NN (total notes) 680 858
CRH rate 27.94% 27.04% LMM 0.006-0.006 z=0.32z=-0.32 0.751
CRNH rate 0.59% 4.55% LMM 0.035-0.035 z=4.13z=-4.13 <0.001<0.001
Helpfulness score 0.31 0.28 LMM $0.019$ $z=2.41$ 0.016
     (NN with scores: 645 / 798)
Num. ratings, median 67.00 91.50 Mann–Whitney $U$ U=236,530.5U=236{,}530.5 <0.001<0.001
Time to CRH, hrs, median 7.16 5.78 Mann–Whitney $U$ U=24,102U=24{,}102 0.098
     (NN CRH notes: 190 / 232)
@@
Metric LLM Human Test AI coef. Statistic $p$
NN (total notes) 102 116
CRH rate 21.57% 20.69% LMM $0.004$ $z=0.09$ 0.929
CRNH rate 1.96% 6.90% LMM 0.049-0.049 z=1.74z=-1.74 0.082
Helpfulness score 0.28 0.24 LMM $0.041$ $z=2.13$ 0.033
     (NN with scores: 96 / 109)
Num. ratings, median 64.00 55.00 Mann–Whitney $U$ U=6,491.5U=6{,}491.5 0.216
Time to CRH, hrs, median 6.01 5.67 Mann–Whitney $U$ $U=288$ 0.605
     (NN CRH notes: 22 / 24)
@@
Metric LLM Human Test AI coef. Statistic $p$
NN (total notes) 190 215
CRH rate 21.58% 22.79% LMM 0.019-0.019 z=0.61z=-0.61 0.542
CRNH rate 2.11% 6.05% LMM 0.039-0.039 z=1.98z=-1.98 0.048
Helpfulness score 0.28 0.25 LMM $0.023$ $z=1.70$ 0.089
     (NN with scores: 178 / 193)
Num. ratings, median 54.00 57.00 Mann–Whitney $U$ U=21,064.5U=21{,}064.5 0.587
Time to CRH, hrs, median 4.61 5.55 Mann–Whitney $U$ $U=950$ 0.662
     (NN CRH notes: 41 / 49)
@@
Metric LLM Human Test AI coef. Statistic $p$
NN (total notes) 257 313
CRH rate 19.46% 20.13% LMM 0.013-0.013 z=0.47z=-0.47 0.638
CRNH rate 1.56% 5.43% LMM 0.037-0.037 z=2.37z=-2.37 0.018
Helpfulness score 0.27 0.24 LMM $0.027$ $z=2.37$ 0.018
     (NN with scores: 240 / 285)
Num. ratings, median 57.00 59.00 Mann–Whitney $U$ U=40,772U=40{,}772 0.778
Time to CRH, hrs, median 4.96 5.46 Mann–Whitney $U$ U=1,562U=1{,}562 0.942
     (NN CRH notes: 50 / 63)

### 13 Apr 2026 v2

@@ 1  Introduction
2  Results
2.1  Platform Dynamics Complicate Note-Level Comparisons
@@
2.2  LLM Notes Receive More Positive Ratings Across the Ideological Spectrum
@@
To examine whether rating patterns varied across the political spectrum, we stratified raters into three ideology groups based on their rater factor: left (factor <0.15<-0.15), neutral ($0.15\leq$ factor $\leq 0.15$), and right (factor >0.15>0.15). For each note and ideology group, we computed the percentage of ratings marked helpful and not helpful, then aggregated across notes to obtain mean percentages with 95% confidence intervals. Results are presented in Figure 1. LLM notes have higher avg. % helpful ratings and lower avg. % unhelpful ratings than human notes across all rater ideology groups. The largest difference appears among neutral raters, followed by left-leaning raters, and right-leaning raters.
To quantify differences in individual ratings between LLM and human notes, we estimated a linear mixed effects model predicting individual rating scores from LLM writer authorship, rater political leaning, and their interactions. This analysis mirrors the “bridging” principle of Community Notes algorithm by modeling how the LLM advantage varies as a function of rater ideology, and the extremity of their ideology. Community Notes allows users to rate each note as Helpful, Somewhat Helpful, or Not Helpful; following the platform convention, we coded these responses numerically as 1.0, 0.5, and 0.0, respectively. Specifically, we use the following primary specification:
$\begin{split}\text{rating\_score}\sim{}&\text{AI}\times\text{coreRaterFactor1}+\text{AI}\times(\text{coreRaterFactor1})^{2}\\ &+(1\mid\text{noteId})+(1\mid\text{raterId})\end{split}$ (1)
@@
Table 1: Rating-Level linear mixed effects model results. To assess robustness, we estimate three additional specifications. Model 2 replaces note random intercepts with tweet random intercepts that absorbs tweet-level confounders while still identifying the AI advantage partly from within-tweet variation. Model 3 is an OLS regression with standard errors clustered by note. Model 4 replaces the continuous coreRaterFactor1 with categorical rater group (left-leaning, neutral, right-leaning; reference: neutral). The estimates of AI advantage stay consistent across all specifications.
Note + Rater RE Tweet + Rater RE OLS Clustered Note + Rater RE (Group) (Intercept) 0.785*** 0.842*** 0.826*** 0.785*** (0.006) (0.005) (0.003) (0.007) AI 0.104*** 0.066*** 0.082*** 0.098*** (0.009) (0.004) (0.004) (0.010) coreRaterFactor1 -0.015*** -0.007 -0.005 (0.004) (0.004) (0.003) coreRaterFactor12 -0.175*** -0.195*** -0.291*** (0.010) (0.010) (0.009) AI:coreRaterFactor1 -0.087*** -0.112*** -0.120*** (0.005) (0.005) (0.005) AI:coreRaterFactor12 -0.190*** -0.169*** -0.181*** (0.013) (0.013) (0.014) left-leaning rater -0.031*** (0.005) right-leaning rater -0.059*** (0.005) AI:left-leaning rater -0.003 (0.007) AI:right-leaning rater -0.086*** (0.007) SD (Intercept raterParticipantId) 0.160 0.160 0.164 SD (Observations) 0.304 0.335 0.304 SD (Intercept noteId) 0.196 0.197 SD (Intercept tweetId) 0.142 Num.Obs. 108169 108169 108169 108169
@@
2.3  Equal-Exposure Note-Level Analysis
@@
2.4  LLM and Human Notes Differ in Writing and Sources
LLM and human notes differed in both writing and source citation behaviors. LLM notes were much longer than human notes (mean 35.8 words vs. 26.9 words, $t=22.190$, $p<0.001$). The LLM writer also cited more URLs on average (mean: 1.51 vs. 1.23 URLs; $t=8.67$, $p<0.001$). Analysis of cited domains reveals distinct sourcing patterns (Tables 2(a) and 2(b)). LLM notes most frequently referenced mainstream news outlets and social media platforms, including reuters.com (7.7% of LLM notes), en.wikipedia.org (7.2%), instagram.com (5.8%), youtube.com (5.3%), x.com (4.2%), bbc.com (4.2%), snopes.com (3.9%), cnn.com (2.4%), facebook.com (2.0%), and yahoo.com (2.0%). In contrast, human notes most commonly cited x.com (18.7%), nearly four times the rate observed in LLM notes (4.2%). Wikipedia appeared at similar rates in both groups (7.6% human vs. 7.2% LLM), but other mainstream news sources appeared less frequently in human notes: reuters.com (1.6% vs. 7.7%), bbc.com (1.1% vs. 4.2%), and snopes.com (0.7% vs. 3.9%). This suggests that the LLM writer relies more heavily on traditional authoritative sources, whereas human writers more frequently reference platform-native content and social media posts. We note that this pattern characterizes our specific implementation.
(a) Top 10 domains in LLM notes.
@@ 2.4  LLM and Human Notes Differ in Writing and Sources
(b) Top 10 domains in human notes.
@@
3  Discussion
@@
4  Data and Methods
4.1  Dataset
@@
4.2  Analysis and multiple testing strategy
@@
4.3  Note Writing Pipeline
@@
Acknowledgements
We thank Dean Eckles, Martin Saveski, Soham De for insightful conversations and feedback on this work. We thank Brad Miller, Jay Baxter, Keith Coleman for valuable discussions on Community Notes and AI fact-checking.
@@ Appendix B Full-Sample Note-Level Outcomes
% Helpful, mean [95% CI] % Unhelpful, mean [95% CI] Rater Ideology LLM Human LLM Human Left ($n$: 1,300 / 1,183) 79.55 [77.83, 81.27] 68.10 [65.95, 70.24] 17.98 [16.33, 19.63] 28.72 [26.61, 30.83] Neutral ($n$: 1,174 / 1,049) 89.58 [88.32, 90.84] 76.83 [74.86, 78.79] 07.60 [6.52, 8.69] 18.40 [16.56, 20.25] Right ($n$: 1,300 / 1,196) 73.13 [71.32, 74.93] 67.46 [65.52, 69.41] 23.65 [21.92, 25.38] 28.58 [26.68, 30.48]
@@
Metric LLM Human Test AI coef. Statistic $p$ NN (total notes) 1,614 1,332 CRH rate 13.07% 18.02% LMM 0.058-0.058 z=4.56z=-4.56 <0.001<0.001 CRNH rate 1.12% 4.13% LMM 0.027-0.027 z=4.71z=-4.71 <0.001<0.001 Helpfulness score 0.25 0.24 LMM $0.007$ $z=1.14$ 0.253      ($N$ with scores: 1,243 / 1,130) Num. ratings, median 22 51 Mann–Whitney $U$ U=740,798U=740{,}798 <0.001<0.001 Time to CRH, hrs, median 7.38 5.90 Mann–Whitney $U$ U=27,979U=27{,}979 0.054      ($N$ CRH notes: 211 / 240)
@@ Appendix C Robustness Checks for Note-Level Outcomes
Metric LLM Human Test AI coef. Statistic $p$ NN (total notes) 680 858 CRH rate 27.94% 27.04% LMM 0.006-0.006 z=0.32z=-0.32 0.751 CRNH rate 0.59% 4.55% LMM 0.035-0.035 z=4.13z=-4.13 <0.001<0.001 Helpfulness score 0.31 0.28 LMM $0.019$ $z=2.41$ 0.016      ($N$ with scores: 645 / 798) Num. ratings, median 67.00 91.50 Mann–Whitney $U$ U=236,530.5U=236{,}530.5 <0.001<0.001 Time to CRH, hrs, median 7.16 5.78 Mann–Whitney $U$ U=24,102U=24{,}102 0.098      ($N$ CRH notes: 190 / 232)
Metric LLM Human Test AI coef. Statistic $p$ NN (total notes) 102 116 CRH rate 21.57% 20.69% LMM $0.004$ $z=0.09$ 0.929 CRNH rate 1.96% 6.90% LMM 0.049-0.049 z=1.74z=-1.74 0.082 Helpfulness score 0.28 0.24 LMM $0.041$ $z=2.13$ 0.033      ($N$ with scores: 96 / 109) Num. ratings, median 64.00 55.00 Mann–Whitney $U$ U=6,491.5U=6{,}491.5 0.216 Time to CRH, hrs, median 6.01 5.67 Mann–Whitney $U$ $U=288$ 0.605      ($N$ CRH notes: 22 / 24)
@@
Metric LLM Human Test AI coef. Statistic $p$ NN (total notes) 190 215 CRH rate 21.58% 22.79% LMM 0.019-0.019 z=0.61z=-0.61 0.542 CRNH rate 2.11% 6.05% LMM 0.039-0.039 z=1.98z=-1.98 0.048 Helpfulness score 0.28 0.25 LMM $0.023$ $z=1.70$ 0.089      ($N$ with scores: 178 / 193) Num. ratings, median 54.00 57.00 Mann–Whitney $U$ U=21,064.5U=21{,}064.5 0.587 Time to CRH, hrs, median 4.61 5.55 Mann–Whitney $U$ $U=950$ 0.662      ($N$ CRH notes: 41 / 49)
@@
Metric LLM Human Test AI coef. Statistic $p$ NN (total notes) 257 313 CRH rate 19.46% 20.13% LMM 0.013-0.013 z=0.47z=-0.47 0.638 CRNH rate 1.56% 5.43% LMM 0.037-0.037 z=2.37z=-2.37 0.018 Helpfulness score 0.27 0.24 LMM $0.027$ $z=2.37$ 0.018      ($N$ with scores: 240 / 285) Num. ratings, median 57.00 59.00 Mann–Whitney $U$ U=40,772U=40{,}772 0.778 Time to CRH, hrs, median 4.96 5.46 Mann–Whitney $U$ U=1,562U=1{,}562 0.942      ($N$ CRH notes: 50 / 63)